If you use 5.76mol of sodium fluoride (NaF) and dissolved this into 3.62 kg of water, what will be the change in the boiling point of your solution. Assume the Kb of water is 0.51°C/m.

1
Expert's answer
2022-03-30T02:33:03-0400

∆Tb= iKbm

m= 5.76÷3.62= 1.5912

i= 2

∆Tb=2×0.51×1.5912

=1.62oc

The boiling point is elevated by 1.62oc